Circular No. 6806

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971227 J. Mendez, Isaac Newton Group (ING), La Palma, and University of

Barcelona (UB); P. Ruiz-Lapuente, UB; and N. Walton, ING, report: “We used the 2.5-m Isaac Newton Telescope at the Observatorio del Roque de los Muchachos to survey the field of GRB 971227, using the Wide Field Camera in the (RGO) I band. We did not detect the GRB afterglow previously reported (IAUC 6800), down to a limiting mag I = 22, ten days after the burst (Jan. 6.1 UT). We note an extended object (possibly a faint galaxy) located 15′′ west and 3′′ south of the position of the GRB reported by Castro-Tirado et al (ibid.). Our images are available at

http://www.ing.iac.es/~jma/grb971227.gif.”

V1333 AQUILAE C. Chevalier and S. A. Ilovaisky, Observatoire de Haute Provence

(OHP), write: “CCD photometry of the optical counterpart of the recurrent x-ray transient Aql X-1 has been obtained with the 1.2-m OHP telescope from 1997 Feb. 19 through Nov. 23. Analysis of the data taken when the source was bright (29 nights during Aug. 6–Sept. 25) shows the presence of the 19-hr modulation reported previously by us (1991, A.Ap. 251, L11). When combined with data taken during the 1995 and 1996 bright states, which also show the modulation, the periodogram of the entire data set (484 frames taken on 59 nights) yields a strong and unique peak with period 0.789 498 ± 0.000 010 day (18.9479 ± 0.0002 hr). An independent time analysis of data taken during quiescence in 1997 (112 frames taken on 27 nights during Mar. 31–July 15 and Oct. 8–Nov. 23) shows a peak at the same value of the period, to within 0.02 percent, as well as a smaller peak at one-half this period. The folded lightcurve is thus double-waved, with unequal minima and slightly unequal maxima. The deepest minimum is in excellent phase agreement with the single bright state lightcurve minimum. We conclude that both the quiescent state and the bright state modulations are directly related to orbital motion.”

COMET 55P/TEMPEL-TUTTLE Visual m_{1} magnitude estimates: Jan. 2.49 UT, 10.8 (C. E. Spratt, Vic-

toria, BC, 0.20-m reflector); 4.19, 10.0 (G. Comello, Roden, The Netherlands, 0.30-m reflector); 8.12, 9.7 (K. Hornoch, Lelekovice, Czech Republic, 0.13-m reflector); 10.43, 9.0 (J. Bortle, Stormville, NY, 20×80 binoculars); 11.30, 9.0 (P. Creed, Massillon, OH, 0.20-m reflector).

1998 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green